OpenStudy (anonymous):

Which lines have negative slopes? Choose all answers that are correct. (Points : 1) x = –3 5x + 2y = 10 –3x – y = 6 –2x + 4y = 16

OpenStudy (igreen):

x = –3 - Undefined slope 5x + 2y = 10 - Negative slope –3x – y = 6 - Negative slope -2x + 4y = 16 - Positive slope You can use this graphing calculator: https://www.desmos.com/calculator

OpenStudy (igreen):

If you see these diamonds: '��' then click out of the post and click back in to get it to normal.